# Multicenter Privacy-Preserving Model Training for Deep Learning Brain Metastases Autosegmentation

### v1

@@ 4 Discussion
In general, F1 is the most widely used Fβ score in medical classification/segmentation tasks. For BM detection, the first primary clinical requirement is to correctly detect the existence of BM with high sensitivity irrespective of the exact 3D extent of each metastasis. False positive BM can be removed afterwards by human expert review. In practice, achieving a high sensitivity is more challenging than achieving a high precision for a deep learning model. For example, a high precision can be easily obtained by a network if it only detects a few large, high-contrast BM (which are easy) without false positives, but it is very challenging to detect all the metastases (large and small, high and low contrast). Therefore, F2, which emphasizes sensitivity more than precision, is more appropriate than F1 as the BM detection metric for clinical application [41]. Fig. 2 (and the average detection performance of Tab. 2 in [29] as well) demonstrates the advantage of F2 over F1 with increasing training data amount. It is worth mentioning that the advantage of LWF over TL and single-center training is more prominent in F1 than F2, as displayed in Tab. 3.
The single-center training shows that Stanford, NYU and BraTS had low F2 scores. With mixed training, the F2 score of Stanford was improved. However, those of NYU and BraTS stayed low despite of improvement. This is very likely because of their distinct image features from other datasets, in particular the low axial slice resolution. This implies that deep learning models for BM identification also require high image quality similar to human experts. For human experts, $\leq$1 mm isotropic resolution on 3D scans has become a consensus of image resolution requirement among radiation oncology communities [31, 30]. Human experts have also reached consensus on optimal sequence selection, contrast agent uptake, distortion correction, and motion control [30]. Satisfying such requirements are believed to improve deep learning model performance for BM identification as well.
@@
Several multicenter studies on deep learning for BM autosegmentation have been published previously [14, 11, 20, 15, 42, 43]. In contrast to the AURORA multicenter study, which reported a median BM volume size of 7.3 cm3, the datasets in our study feature significantly smaller median volume sizes, all less than 0.5 cm3. The principal novelty of our work lies in the introduction of a privacy-preserving method, i.e. LWF, which enables the training of a model across multiple centers without the need to share raw data. This contrasts with other studies, which typically rely on pooled multicenter data for model training. This approach addresses significant privacy concerns and enhances the feasibility of multicenter collaboration on BM autosegmentation.
The proposed LWF method generally exhibits robustness to parameter selection. The parameter $\lambda$ in Eqn. 1 governs the balance between stability and plasticity [24]. A high $\lambda$ value results in minimal model alterations at the target center, maintaining stability, whereas a low $\lambda$ value aligns the model’s performance more closely with that achieved by naive TL. Our experiments suggest that a $\lambda$ range from 0.01 to 0.5 effectively balances stability and plasticity, with 0.1 identified as the optimal value. To facilitate effective multicenter collaboration, it is crucial that datasets from different centers undergo the same preprocessing steps, as detailed in Section 2.1. This standardization ensures consistency across centers, enhancing the overall efficacy of the collaboration and ultimately improving model generalizability.
@@
In addition to further combinations of datasets in this work, future studies will aim to include a wider variety of datasets from different hospitals and institutions across Europe. This would test the model’s robustness and ability to generalize across diverse patient demographics and equipment variations. The ultimate goal is to seamlessly integrate the method in this work into existing clinical workflows. Achieving this integration necessitates comprehensive AI-specific quality assurance (QA) measures [44] to ensure the reliability, accuracy, and safety of the model. Prior to clinical deployment, the model must undergo stringent validation and secure approval from relevant health authorities, such as the Food and Drug Administration (FDA) or European Medicines Agency (EMA), confirming its safety and efficacy within clinical environments. Significant additional research will be necessary to bridge the gap between the developmental phase of this deep learning model for BM autosegmentation and its practical clinical application, ensuring that all scientific, regulatory, and operational benchmarks are met.

### v2

@@ 4 Discussion
Several multicenter studies on deep learning for BM autosegmentation have been published previously [10, 8, 9, 11, 12, 13]. In contrast to the AURORA multicenter study with a median BM volume size of 7.3 cm3, the datasets in our study feature significantly smaller median volume sizes, all less than 0.5 cm3. The principal novelty of our work lies in the introduction of a privacy-preserving method, i.e. LWF, which enables the training of a model across multiple centers without the need to share raw data. This contrasts with other studies, which typically rely on pooled multicenter data for model training. This approach addresses significant privacy concerns and enhances the feasibility of multicenter collaboration on BM autosegmentation.
The single-center training shows that Stanford, NYU and BraTS had low F2 scores. With mixed training, the F2 scores of NYU and BraTS stayed low despite of improvement. This is very likely because of their distinct image features from other datasets, in particular the low axial slice resolution. This implies that deep learning models for BM identification also require high image quality similar to human experts. For human experts, $\leq$1 mm isotropic resolution on 3D scans has become a consensus of image resolution requirement among radiation oncology communities [23]. Human experts have also reached consensus on optimal sequence selection, contrast agent uptake, distortion correction, and motion control [23]. Satisfying such requirements are believed to improve deep learning model performance for BM identification as well.
In general medical tasks, F1 is the most widely used Fβ score. For BM detection, the first primal clinical requirement is high sensitivity in detecting BM, regardless of the exact 3D extent of each metastasis. False positives can be removed afterwards by human expert review. In practice, achieving high sensitivity is more challenging than achieving high precision. For example, a network can achieve high precision by detecting a few large, high-contrast BM, but it is very challenging to detect all the metastases (large and small, high and low contrast). Therefore, F2, which prioritizes sensitivity, is more appropriate than F1 for BM detection in clinical application. Fig. 2(a)-(b) and the average detection performance of Tab. 2 in [21] demonstrate the advantage of F2 over F1 with increasing training data amount. Note that the benefit of LWF over TL and single-center training is more prominent in F1 than F2, as displayed in Fig. 2 and Tab. 3.
@@
In addition, future studies will aim to include a wider variety of datasets from different hospitals and institutions across Europe. This would test the model’s robustness and ability to generalize across diverse patient demographics and equipment variations. The ultimate goal is to seamlessly integrate the method in this work into existing clinical workflows. Achieving this integration necessitates comprehensive AI-specific quality assurance (QA) measures [36] to ensure the reliability, accuracy, and safety of the model. Prior to clinical deployment, the model must undergo stringent validation and secure approval from relevant health authorities, such as the Food and Drug Administration (FDA) or European Medicines Agency (EMA). Significant additional research will be necessary to bridge the gap between the developmental phase and practical clinical applications, ensuring that all scientific, regulatory, and operational benchmarks are met.